OpenStudy (anonymous):

solve by factoring. 4x^2+23x+15=0

OpenStudy (nathan917):

Simplifying 4x2 + 23x + 15 = 0 Reorder the terms: 15 + 23x + 4x2 = 0 Solving 15 + 23x + 4x2 = 0 Solving for variable 'x'. Factor a trinomial. (5 + x)(3 + 4x) = 0 Subproblem 1 Set the factor '(5 + x)' equal to zero and attempt to solve: Simplifying 5 + x = 0 Solving 5 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + x = 0 + -5 Combine like terms: 5 + -5 = 0 0 + x = 0 + -5 x = 0 + -5 Combine like terms: 0 + -5 = -5 x = -5 Simplifying x = -5 Subproblem 2 Set the factor '(3 + 4x)' equal to zero and attempt to solve: Simplifying 3 + 4x = 0 Solving 3 + 4x = 0 Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + 4x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + 4x = 0 + -3 4x = 0 + -3 Combine like terms: 0 + -3 = -3 4x = -3 Divide each side by '4'. x = -0.75 Simplifying x = -0.75 Solution x = {-5, -0.75} forgive me was afk :P

OpenStudy (anonymous):

first note, everything is positive- no negatives when factoring next, look at multiple pairs of 4 and 15 mulitple pairs 4- (1 4) (2 2) 15- (1 15) (3 5) the determine if a combination/ dot product from 4 and 15 will give you 23 1*1+4*15 = 61 1*15+4*1=19 2*1+2*15=32 etc once you find the combination, then arrange in the form of (ax+b)(cx+d)=0 where a and c were the multiple pairs from 4 and b and d were the multiple pairs from 15

OpenStudy (anonymous):

thanks guys (: